OpenStudy (anonymous):

Of 1232 people who came into a blood bank to give blood, 397 people had high blood pressure. Estimate the probability that the next person who comes in to give blood will have high blood pressure

OpenStudy (anonymous):

the ratio \(\frac{397}{1232}\)

OpenStudy (anonymous):

which, if you were going to estimate, would be \(\frac{400}{1200}=\frac{1}{3}\)

OpenStudy (anonymous):

it has to be a decimal

OpenStudy (anonymous):

why? one third makes a great fraction, and a lousy decimal

OpenStudy (anonymous):

0.241 0.29 0.322 0.373 those are my answer choices